# RedDebate: Safer Responses through Multi-Agent Red Teaming Debates

### v1

@@ 4.2 Evaluation Metrics
For each agent $n$, given the evaluated safety label $y_{p,n}^{(t)}$ at round $t$ for the $p$-th input prompt, we assess the safety of debates using two metrics:
@@ 4.3 Implementation Details
By default, all debates are conducted over three rounds. We use Mistral-7B-v0.2, LLaMA-3.2-3B, and Phi-3.5-mini as our primary backbone models for the debater agents. In the Devil–Angel setting, we use these three models and explored all possible role assignment, reporting the average performance across permutations. For the feedback generator and Socratic agent, which play critical guiding roles, we use GPT-4o-mini to mitigate context window limitations. All safety evaluations are performed with LlamaGuard. The complete agent prompts are listed in Appendix I.
@@ 5.1 Debate Performance
Scenario HarmBench CoSafe
Error Rate (%) $\downarrow$ Agreement Rate (%) $\uparrow$ Error Rate (%) $\downarrow$ Agreement Rate (%) $\uparrow$
@@
Scenario HarmBench Error Rate (%) $\downarrow$ CoSafe Error Rate (%) $\downarrow$
@@
Engaging in Debate Leads to Safer Answers.
As shown in Table 1, Standard Prompting—where agents independently answer in a single round—exhibits the highest error rate. In contrast, PReD significantly reduces both total and individual error rates, confirming that multi-agent interaction improves response safety. Refinement is a key aspect which captures the improvement in response safety throughout the course of a debate, quantified by reductions in agents’ error rates and increases in agreement. When exploring the impact of different debate scenarios on refinement, SReD achieves the highest agreement rate (17.0% in HarmBench and 3.8% in CoSafe) and the least error rate. This indicates that agents revise unsafe responses more effectively when prompted by explicit questioning or counterarguments from persona agents. Indeed, the more actively agents are engaged in the debate process, the more opportunities they have to refine unsafe reasoning.
Learning from Experience Lowers Error Rates.
We select the best-performing debate scenario without memory—SReD—and equip debaters with different types of LTM to determine whether agents being able to leverage previously learned experiences reduces error rates. As shown in Table 2, when utilizing LTM we observe consistent improvements in overall error rate across all LTM types. Additionally, combining TLTM and CLTM into a Unified LTM leads to even greater error reduction, confirming the benefit of integrating both mechanisms. This highlights that both memory types on their own are not fully sufficient for imparting the information from the generated feedback into the debaters. By integrating both mechanisms a form of constructive interference is created and the information is better relayed allowing debaters to better recount crucial feedback. An ablation on PReD with LTM is also presented in Appendix D, demonstrating that LTM remains effective even in the absence of auxiliary agents. Furthermore, as shown in Appendix E, memory-augmented agents maintain their helpfulness. This is reflected in both accuracy remaining at similar levels and a low refusal rate on safe general questions, which demonstrates that the safety enhancements provided by LTM do not diminish the agents’ overall utility.
Diverse Peer Feedback with Memory Outperforms Self-Critique.
Our method enables agents to receive dynamic feedback, without requiring human input, that is directly grounded in their own mistakes—resulting in more targeted refinement. This allows it to outperform settings where agents refine themselves using fixed, predefined rules, as explored in the Self-Critique baseline (Table 2).
The addition of LTM significantly enhances debate performance relative to Self-Critique, despite Self-Critique initially achieving lower error rates on HarmBench compared to debate scenarios without memory (e.g., 15.36% vs. 21.0% in SReD). Self-Critique likely has an initial edge due to the simplicity and the effect of human input—agents are simply given their own previous response and asked to revise it per a predefined constitutional rule. In contrast, debate agents must reason and respond to a dynamic conversation without direct guidance which adds greater complexity. Hence, LTM allows for additional guidance for the debaters through feedback while benefiting from the additional depth and reasoning provided from the debate process, allowing for optimal performance.
@@ 5.2 Detailed Analysis
As shown in Table 2, the “SReD+gltm” configuration yields the lowest error rates across both HarmBench and CoSafe, demonstrating the value of converting learned unsafe behaviours into explicit control flows that proactively block harmful inputs. Its effectiveness stems from the fact that “prevention is better than cure”: harmful inputs that match prior unsafe patterns are intercepted before reaching the model. In addition, Colang’s intent detection follows a two-stage process—similar to RankGPT (Sun et al., 2023)—where it first retrieves relevant guardrails and then uses a generative agent to assess whether any match is strong enough to trigger rejection. This layered verification offers greater precision than relying solely on retrieval or embedding similarity.
Among the models, Llama shows the most reliable GLTM integration, with the lowest error rates (0.3% on HarmBench, 0.4% on CoSafe), high recall of unsafe intents (99.4%, 92.7%), and minimal runtime failures, making it well-suited for guardrail-based safety (Table 3). In contrast, Phi suffers from non-negligible runtime errors due to NeMo’s instability and formatting constraints.
Metric HarmBench CoSafe
Mistral Llama Phi Mistral Llama Phi
Intent Match (%) $\uparrow$ 25.5 39.5 26.0 2.4 4.5 1.4
Guardrails Recall (%) $\uparrow$ 76.7 99.4 94.5 35.4 92.7 35.2
Runtime Error (%) $\downarrow$ 7.3 0.3 25.0 0.0 0.1 9.7
Category-Level Errors.
We analyze category-level error distributions to identify areas where agents are more prone to unsafe responses. SReD consistently performs better across most categories; however, challenges persist in categories such as copyright and chemical biological in HarmBench, and terrorism and financial harm in CoSafe. For more details and visualizations, see Appendix F.
(a) Debate
(b) Self-Critique
Figure 3: Stepwise error rates for Debate and Self-Critique. Debate shows consistent error reduction and convergence, while Self-Critique improves less reliably.
Refinement Dynamics.
To closely examine the refinement dynamics of debate and Self-Critique, we analyze step-wise error rates (Figure 3). Debaters with memory consistently reduce errors across each round (Figure 3(a)), eventually converging toward a similar error rate range. This convergence suggests that debaters achieve a safe consensus through interactive debate. Conversely, Self-Critique (Figure 3(b)) shows no convergence due to isolated revisions, highlighting debate’s advantage: external perspectives and peer feedback are crucial for effective safety refinement. Agents often struggle to recognize their own unsafe behaviour, which can instead be surfaced and corrected through peer interaction during debate.
@@
We observe that introducing varied debate strategies serves as a mechanism to reveal hidden vulnerabilities not visible in single-turn settings. A sample of this behaviour is shown in Appendix H.2. Taking HarmBench as a case study, we note that Llama and Phi, despite initially providing safe responses (Figure 4(a)), produce answers that become unsafe as the debate progresses (Figure 4(b)). Often, these initial responses are minimal and cautious—such as “I cannot provide you with the answer”—but as agents engage further in the debate or address questions raised by other agents, they fail to maintain safety and ultimately produce unsafe responses. Indeed, this process of error discovery aligns well with the dual focus of our framework on learning and red-teaming. By surfacing previously unexposed unsafe spots, agents are given the opportunity to learn from accompanying feedback and apply those lessons in similar future situations.
(a) Initial Unsafe Responses
@@
Figure 4: Vulnerability heatmaps showing model unsafety on the first attempt and subsequent ones
[…]
(b) Average Steps to Unsafe
Figure 5: Comparison of attack effectiveness (left) and efficiency (right) across approaches.
@@
The rate at which agents transition from safe to unsafe responses depends on the debate strategy. As illustrated in Figure 5(a), the DAReD scenario triggers the highest number of these conversions, whereas SReD results in fewer conversions overall. However, for the cases where SReD does discover vulnerabilities, it uncovers them in fewer average steps compared to other scenarios. To further analyze the efficiency of debate strategies as “attack methods” that trigger unsafe responses, we plot the results across different models and debate scenarios in Appendix, Figure 6.
Ablations.
Ablation studies on extending debate rounds up to five, using additional peer debaters, and evaluating the effectiveness of evaluator agent are presented in Appendix B. Overall, we show that three debate rounds lead to optimal performance, using additional agents can improve error rate and diversity, and that LlamaGuard produces strong evaluations for safety.
@@ 6 Conclusion
Through our RedDebate framework, we show that multi-agent debates combined with automated red-teaming and long-term memory can significantly enhance LLM safety without human supervision. By enabling agents to critique and refine each other’s responses, our framework both effectively reduces unsafe outputs and uncovers vulnerabilities. Memory modules and proactive guardrails further amplify safety improvements, demonstrating that structured collaboration and systematic feedback offer a practical path toward more robust and scalable AI safety solutions.
@@
As with any automatic tool for evaluation, our primary safety evaluator LlamaGuard does not achieve perfect performance, which can cause some issues in evaluating responses and generating the most optimal feedback. Overall, it does not compromise the performance of our framework and we observe strong increases in safety particularly on HarmBench which cannot be explained by noise in the evaluator. Additionally, we cover very standard benchmark datasets in HarmBench and CoSafe which span a wide range of domains, although we do not focus on underrepresented domains in this work as we aim to achieve a broad view of the issue.
We focus our experiments on open-source, small scale language models, as these are more likely to suffer from misalignment from human values and stand to gain the most improvement in safety through our approach, which we demonstrate through our results. These models are especially useful because of their transparency and ability to be applied in domains where API-based models are unsuitable. It is possible to apply our framework to large state-of-the-art language models offered through API services, but these models undergo far more alignment and extensive patching to become more safe OpenAI (2024), and hence we prioritize testing our framework on the aforementioned models.
A few limitations stem from strict Colang formatting and some of NeMo’s instability led to runtime errors in some models such as Phi, indicating that the library itself needs improvements. Lastly, despite our RedDebate framework leading to a significant reduction in unsafe behaviours, we note that there is an increased inference cost when using our standard debate settings with three agents across three rounds which can increase wall-clock inference by roughly 3-4 times over single-turn baselines. Nevertheless, we stress that this framework aims to maximize safety in applications where it is the upmost concern over fast computation.
@@
This research investigates methods to enhance the safety of large language models through automated debate and red-teaming, with the goal of proactively mitigating harmful or unsafe outputs. All experiments were conducted using publicly available models and datasets. The primary datasets, HarmBench and CoSafe, contain adversarial prompts focused on conversational and social safety; care was taken to ensure that outputs and analysis remained within ethical and legal guidelines. No private or sensitive user data was used.
While our framework aims to reduce unsafe behaviours, automated safety evaluation and guardrails are not foolproof and may miss nuanced or context-dependent harms. We caution that models—even when improved by our techniques—should not be deployed in high-stakes or real-world scenarios without thorough human oversight and external auditing. Our code and results are shared to promote transparency and reproducibility.
The potential misuse of automated debate systems for adversarial or malicious purposes is a recognized risk. We encourage responsible research practices and urge practitioners to consider societal impacts, bias propagation, and unintended consequences when building on or deploying similar methods.
No human subjects, personally identifiable information, or sensitive data were involved in this research.

### v2

@@ 4.2 Evaluation Metrics
For each agent $n$, given the evaluated safety label $y_{p,n}^{(t)}$ at round $t$ for the $p$-th input red-teaming prompt, we assess debates effectiveness using two metrics:
@@ 4.3 Implementation Details
All debates are conducted over three rounds. We use two triads of debater agents: (Mistral-7B-v0.2, LLaMA-3.2-3B-Instruct, Phi-3.5-mini) and (Gemma-3-12B, Qwen3-8B, Deepseek-R1-Distill-LLaMA-8B). In the Devil–Angel setting, we assign roles across all permutations of the three models and report the average performance. For the feedback generator and Socratic agent, which play key guiding roles, we employ GPT-4o-mini to handle context window limitations while leveraging its strong safety performance. All safety evaluations are conducted with LlamaGuard. Complete agent prompts are provided in Appendix F.
@@ 5.1 Debate Performance
Method HarmBench CoSafe
Error Rate (%) ↓ Agreement Rate (%) ↑ Error Rate (%) ↓ Agreement Rate (%) ↑
@@
Total Gemma Qwen R1 Total Gemma Qwen R1 Total Gemma Qwen R1 Total Gemma Qwen R1
Std. Prompting 44.4 47.3 41.3 44.8 7.9 11.8 7.0 5.5
PReD 34.7 29.0 35.1 40.1 13.5 15.0 12.0 13.6 6.8 9.8 5.6 4.8 3.1 4.2 2.8 2.2
DAReD 36.4 31.8 33.4 43.9 14.0 19.4 12.0 10.7 6.9 9.7 5.9 5.2 3.6 5.3 3.3 2.1
SReD 28.9 24.8 28.6 33.4 17.5 20.5 15.8 16.1 5.8 7.9 4.9 4.5 3.8 5.6 3.4 2.4
Table 1: Error rates and Agreement rates (%) across different scenarios (without LTM).
Engaging in Debate Leads to Safer Answers.
As shown in Table 1, Standard Prompting, where agents independently answer in a single round, exhibits the highest error rate. In contrast, PReD significantly reduces both total and individual error rates, confirming that multi-agent interaction improves response safety. Refinement is a key aspect which captures the improvement in response safety throughout the course of a debate, quantified by reductions in agents’ error rates and increases in agreement rates. When exploring the impact of different debate scenarios on refinement, SReD achieves the highest agreement rate and the lowest error rate. This indicates that agents revise unsafe responses more effectively when prompted by explicit questioning or counterarguments from persona agents. Indeed, the more actively agents are engaged in the debate process, the more opportunities they have to refine unsafe reasoning.
Learning from Experience Lowers Error Rates.
We select the best-performing debate scenario without memory—SReD—and equip debaters with different types of LTM to determine whether agents being able to leverage previously learned experiences reduces error rates. As shown in Table 2, when utilizing LTM we observe consistent improvements in overall error rate across all LTM types. Additionally, combining TLTM and CLTM into a Unified LTM in most cases leads to even greater error reduction, confirming the benefit of integrating both mechanisms. This highlights that both memory types on their own are not fully sufficient for imparting the generated feedback into the debaters. Integrating both mechanisms creates a form of constructive interference, which helps relay the information more effectively and enables debaters to better recount crucial feedback. An ablation on PReD with LTM is also presented in Appendix B.6, demonstrating that LTM remains effective even in the absence of auxiliary agents. While LTM enhances safety, it is important to ensure that memory-augmented agents retain their helpfulness. As shown in Appendix B.4, on safe general questions, accuracy remains comparable and refusal rates stay low, demonstrating that LTM’s safety improvements do not compromise overall utility.
Diverse Peer Feedback with Memory Outperforms Self-Critique.
The Self-Critique baseline (Table 2) follows the Constitutional AI framework (Bai et al., 2022b), in which models iteratively revise their responses based on a set of predefined, human-crafted rules (constitutions). This revision process can be repeated over multiple iterations, allowing the model to reflect on its outputs using randomly selected constitutions. In contrast, our method supplies agents with dynamic peer feedback grounded in their own mistakes, requiring no human input. This enables more targeted refinement and outperforms Self-Critique with the same number of revision steps (see Figure 8 in the Appendix). The addition of LTM significantly enhances debate performance relative to Self-Critique, despite Self-Critique initially achieving lower error rates on HarmBench compared to debate scenarios without memory (e.g., 15.36% vs. 21.0% in SReD for the first triad). Self-Critique likely has an initial edge due to the simplicity and the effect of human input—agents are simply given their own previous response and asked to revise it per a predefined rule. In contrast, debate agents must reason and respond to a dynamic conversation without direct guidance which adds greater complexity. Hence, LTM provides the necessary guidance while leveraging the added depth and reasoning of the debate process, allowing for optimal performance.
Scenario HarmBench Error Rate (%) ↓ CoSafe Error Rate (%) ↓
@@
Total Gemma Qwen R1 Total Gemma Qwen R1
Self-Critique 8.3 8.1 8.5 8.3 11.2 10.6 9.9 13.1
SReD 28.9 24.8 28.6 33.4 5.8 7.9 4.9 4.5
+tltm 6.4/+22.5 4.5/+20.3 6.6/+22.0 8.0/+25.4 3.4/+2.4 2.1/+5.8 1.2/+3.7 6.9/-2.4
+cltm 12.6/+16.3 5.1/+19.7 2.6/+26.0 30.1/+3.3 4.9/+0.9 3.7/+4.2 3.7/+1.2 7.4/-2.9
+tltm+cltm 2.3/+26.6 0.0/+24.8 1.4/+27.2 5.5/+27.9 2.8/+3.0 1.4/+6.5 4.2/+0.7 2.8/+1.7
+gltm 0.2/+28.7 0.3/+24.50 0.0/+28.6 0.3/+33.1 0.8/+5.0 1.2/+6.70 0.2/+4.7 1.1/+3.4
@@
Debate is Not Just a Compute Hack.
Prior work has shown that scaling up inference-time compute generally improves robustness Zaremba et al. (2025); Wu et al. (2025). Debate, which involves multi-agent reasoning and iterative discussion over a prompt, follows a similar trajectory, but with important additional benefits beyond mere scaling. Unlike a single agent revising in isolation, debate leverages the diversity of perspectives across agents, enabling a more robust safety inference-time scaling process. To investigate this distinction, we analyze the step-wise refinement behaviour of self-critique and debater agents under equivalent inference calls (Figure 3). Debaters with memory steadily reduce error rates over successive rounds, achieving lower step-wise errors. In contrast, self-critique revisions remain isolated and lack external correction. Notably, discrepancies among agents, which are inherent to debate and absent in self-critique, drive safer outcomes by enabling more effective refinement and inference-time scaling; see Appendix B.3 for a detailed analysis.
Figure 3: Stepwise error rates for debate and self-critique.
Figure 4: Self-critique under extended revisions with matched token budget versus debate.
Additional Compute for Self-Critique Does Not Match Debate’s Safety Gains.
We observe that the interactive nature of debate, where agents pose and respond to each other’s arguments, naturally leads to increased token generation. To ensure a fair comparison across methods, we capped each response at $512$ newly generated tokens, preventing debater agents from disproportionately expanding their computational budget. Despite this constraint, self-critique agents produce approximately 35% fewer tokens per turn than debaters ($1,351$ vs. $2,104\pm 228$ tokens on average). We argue that this modest increase in compute is justified by the substantially greater safety gains achieved through debate, gains that self-critique alone fails to replicate. To probe this further, we increase the revision steps for self-critique agents to match the total tokens generated by debaters, although these extra inference calls introduce additional overhead and latency compared to debate. As shown in Figure 4, even with additional revision steps under an equivalent output token budget, self-critique still underperforms relative to debate with just three turns. This finding reveals a key insight that isolated agents, even when granted more compute, often fail to identify and correct their own unsafe blind spots. In contrast, the interactive exchange in debate facilitates more effective and robust error correction. While debate incurs slightly higher per-turn costs, its substantial benefits in improving safety make it a compelling approach, particularly in settings where safety is paramount.
@@ 5.2 Detailed Analysis
As shown in Table 2, the “SReD+gltm” configuration yields the lowest total error rates across HarmBench and competitive performance on CoSafe, demonstrating the value of converting learned unsafe behaviours into explicit control flows that proactively block harmful inputs. Its effectiveness stems from the fact that “prevention is better than cure”: harmful inputs that match prior unsafe patterns are intercepted before reaching the model. In addition, Colang’s intent detection follows a two-stage process, similar to RankGPT (Sun et al., 2023), where it first retrieves relevant guardrails and then uses a generative agent to assess whether any match is strong enough to trigger rejection. This layered verification offers greater precision than relying solely on retrieval or finetuning. However, for some models such as Phi, GLTM can often cause some non-negligible runtime errors due to NeMo’s instability and formatting constraints. See Appendix C.3 for details.
@@
We observe that introducing varied debate strategies serves as a mechanism to reveal hidden vulnerabilities not visible in single-turn settings. A sample of this behaviour is shown in Appendix E.2. Taking HarmBench and the first triad as a case study, we note that Llama and Phi, despite initially providing safe responses (Figure 5(a)), produce answers that become unsafe as the debate progresses (Figure 5(b)). Often, these initial responses are minimal and cautious, such as “I cannot provide you with the answer”, but as agents engage further in the debate or address questions raised by other agents, they fail to maintain safety and ultimately produce unsafe responses. Indeed, by revealing previously unexposed unsafe spots, error discovery allows agents to learn from the accompanying feedback and apply these lessons in similar situations.
(a) Initial Unsafe
@@
(c) Conversion Rate
(d) Steps to Unsafe
Figure 5: Vulnerability heatmaps and attack metrics across settings.
@@
The rate at which agents transition from safe to unsafe responses depends on the debate strategy. As illustrated in Figure 5(c), the DAReD scenario triggers the highest number of these conversions, whereas SReD results in fewer conversions overall. However, for the cases where SReD does discover vulnerabilities, it uncovers them in fewer average steps compared to other scenarios (Figure 5(d)). To further analyze the efficiency of debate strategies as “attack methods” that trigger unsafe responses, we plot the results across different models and debate scenarios in Appendix, Figure 12. We also analyze category-level error distributions to identify areas where agents are more prone to unsafe responses in Appendix B.9.
Validating LLM-Based Safety Evaluators Against Human Judgment
Relying solely on LLM-based evaluators for safety assessment can perpetuate biases and miss nuanced human judgments, potentially obscuring misalignment Panickssery et al. (2025). To address this, we evaluated LlamaGuard against human-labeled safety judgments on a subset of generated debate arguments. LlamaGuard performed reliably, supporting its use as a proxy for human evaluation in our setting. This validation provides confidence that safety improvements on benchmarks reflect alignment with human values while still benefiting from the automation offered by the framework (see Appendix B.2).
Ablation Studies on Debater Behavior
We examine the behaviour of strongly aligned models such as GPT-OSS in adversarial debates and find that, while they avoid unsafe outputs, they rarely engage with peers on red-teaming prompts, limiting their effectiveness as responsible agents in multi-agent discussions (see results in Appendix B.1). We also study the impact of instruction-tuning by replacing Llama with its base version, observing that the base model fails to participate meaningfully in debate, causing unsafe content to persist more than in the tuned model (Appendix B.5).
Ablation Studies on Debate Rounds and Peer Debaters
We conduct ablation studies on extending the number of debate rounds and peer debaters, each up to five (see Appendix B.7 and Appendix B.8). Our findings indicate that three debate rounds yield optimal performance, and that incorporating additional debater agents improves both error reduction and response diversity.
@@ 6 Conclusion
We introduce the RedDebate framework and show that multi-agent debates combined with automated red-teaming and long-term memory can significantly enhance LLM safety without human supervision. By enabling agents to critique and refine each other’s responses, our framework both effectively reduces unsafe outputs and uncovers vulnerabilities. Memory modules and proactive guardrails further amplify safety improvements, demonstrating that structured collaboration and systematic feedback offer a practical path toward more robust and scalable AI safety solutions.
@@
As with any automatic evaluation tool, our primary safety evaluator LlamaGuard, while performing acceptably compared to human baseline annotations (as shown in our ablation studies), does not achieve perfect performance. This can cause some issues in evaluating responses and generating the most optimal feedback. Overall, it does not compromise the performance of our framework and we observe strong increases in safety particularly on HarmBench which cannot be explained by noise in the evaluator. Additionally, we cover very standard benchmark datasets in HarmBench and CoSafe which span a wide range of domains, although we do not focus on underrepresented domains in this work as we aim to achieve a broad view of the issue.
We focus our experiments on open-source, small scale language models, as these are more likely to suffer from misalignment from human values and stand to gain the most improvement in safety through our approach, which we demonstrate through our results. These models are especially useful because of their transparency and ability to be applied in domains where API-based models are unsuitable. It is possible to apply our framework to large state-of-the-art language models offered through API services, but these models undergo far more alignment and extensive patching to become more safe OpenAI (2024). They also frequently use external methods and modules to handle unsafe content making it difficult to determine the model’s true capabilities, and hence we prioritize testing our framework on the aforementioned models.
A few limitations stem from strict Colang formatting and some of NeMo’s instability led to runtime errors in some models such as Phi, indicating that the library itself needs improvements.
Finally, although this approach requires more computational resources per step, we maintain that the substantial safety improvements justify this investment. While the additional inference time occurs during training and adaptation phases rather than deployment, we emphasize that this framework is specifically designed for applications where safety takes precedence over computational speed. Once models have participated in debate and their learned lessons have been distilled into long-term memory, the cost to inference is not heavy. We also believe that it is crucial to introduce and validate these safety-focused approaches early, as ongoing improvements in computational hardware are expected to make inference more affordable, thereby enhancing the practical viability of safety-focused methodologies like debate.
Use of Large Language Models
All content, analyses, and results in this paper are written in our own writing style; LLMs were only used for minor grammar checks and word choice refinements after initial draft, ensuring clarity and readability without altering the scientific content or original presentation of ideas.
@@
This research investigates methods to enhance the safety of large language models through automated debate and red-teaming, with the goal of proactively mitigating harmful or unsafe outputs. All experiments were conducted using publicly available models and datasets. The primary datasets, HarmBench and CoSafe, contain adversarial prompts focused on conversational and social safety; care was taken to ensure that outputs and analysis remained within ethical and legal guidelines. No private or sensitive user data was used. While our framework aims to reduce unsafe behaviours, automated safety evaluation and guardrails are not foolproof and may miss nuanced or context-dependent harms. We caution that models—even when improved by our techniques—should not be deployed in high-stakes or real-world scenarios without thorough human oversight and external auditing. Our code and results are shared to promote transparency and reproducibility. The potential misuse of automated debate systems for adversarial or malicious purposes is a recognized risk. We encourage responsible research practices and urge practitioners to consider societal impacts, bias propagation, and unintended consequences when building on or deploying similar methods. No human subjects, personally identifiable information, or sensitive data were involved in this research.